# The equivalence of asymptotic average shadowing and vague specification properties and its consequences

### v1

@@ 6 Examples of shift spaces with the vague specification property
6.1 An example of proximal shift space
We will show that the shift space $Z$, defined in Example 21 in [6], has the vague specification property. Since it is known that $Z$ has the $\bar{d}$-shadowing property, it is enough to show that $Z$ is $\bar{d}$-complete and use Corollary 5.17. Before showing that $Z$ is $\bar{d}$-complete we recall its construction following [6].
@@ 6.1 An example of proximal shift space
We let $\tilde{E}_{j}$ and $E_{i}$ to be defined as in (9). Let $\{x^{(n)}\}_{n=1}^{\infty}$ be a $\bar{d}$-Cauchy sequence in $Z$. Passing to a subsequence, we can assume that for every $n\in\mathbb{N}$ it holds
@@
Note that $Y_{n}$ is a hereditary shift with specification for every $n\in\mathbb{N}$. Therefore, by Lemmas 5.1 and 5.2, for every $n\in\mathbb{N}$ there exists $m_{n}\in\mathbb{N}$ such that $u_{n}=0^{m_{n}}$ is a synchronizing word for $Y_{n}$.
@@
Note that the definition of $Z$ implies that $y^{(n)}\in Z$. Observe that, using (10), for every $n\in\mathbb{N}$ we have
@@
The proximal shift space $Z$ has the asymptotic average shadowing property.
6.2 A class of minimal shift spaces
In [6], the authors provided a way of constructing minimal shift spaces that are mixing and have the $\bar{d}$-shadowing property, acknowledging Piotr Oprocha for the ideas that inspired this construction. It is worth it mentioning that the obtained shift spaces have positive topological entropy (for definition and details we refer [6]). We will show that those shift spaces are $\bar{d}$-complete.
@@ 6.2 A class of minimal shift spaces
For $n\in\mathbb{N}$, let $X_{n}$ be the coded shift space generated by the code ${\mathcal{B}}_{n}$. That is, $X_{n}$ consists of all infinite concatenations of words from ${\mathcal{B}}_{n}$ together with their shifts. The proofs of the facts below can be found in Section 4.2 of [6].
@@
For every $n\in\mathbb{N}$ it holds $X_{n+1}\subseteq X_{n}$, which implies $\bigcap_{n=1}^{\infty}X_{n}\neq\emptyset$.
@@
From now on, we set ${\mathcal{B}}_{1}=\{0,11\}$. As noted in [6], from this choice of ${\mathcal{B}}_{1}$, a simple inductive argument shows that, for every $n\in\mathbb{N}$, the set of lengths of all words in ${\mathcal{B}}_{n}$ is an interval.
@@
The next lemma is a direct consequence of Proposition 6.8 and was extracted from the proof of Proposition 25 in [6]. For the sake of completeness we will replicate the argument.
@@
In [6], it is proved that under certain conditions on the sequence $\{t(n)\}_{n=1}^{\infty}\subset\mathbb{N}$ the obtained shift space $X$ possess various dynamical properties.
Theorem 6.10 (Theorem 27 in [6]).
@@
The sequences $\{t(n)\}_{n=1}^{\infty}$ mentioned on Theorem 6.10 fulfill the following conditions for every $n\in\mathbb{N}$:
@@
From now on, we will assume that the sequence $\{t(n)\}_{n=1}^{\infty}$ fulfills the conditions above. We will show that these under these conditions the obtained shift space $X$ is $\bar{d}$-complete.
@@
We can improve Lemma 6.11 to ensure that the prefix of the obtained word $w$ can be chosen, while also providing more precise estimates for the number of distinct letters from $x_{[0,|w|)}$ and $w$ when analyzing subwords.
@@
Let $\{x^{(n)}\}_{n=1}^{\infty}\subset X$ be a $\bar{d}$-Cauchy sequence. Up to pass to a subsequence we can assume that for every $n\in\mathbb{N}$ it holds
@@
We can construct a strictly increasing sequence $\{k_{n}\}_{n=1}^{\infty}\subset\mathbb{N}$ such that for every $n\in\mathbb{N}$ and $\ell\geq k_{n}$ it holds
@@
We will construct a strictly increasing sequence $\{m_{j}\}_{j=1}^{\infty}\subset\mathbb{N}$ and a sequence of words $\{w^{(j)}\}_{j=1}^{\infty}\subset\mathcal{L}(X)$ such that for all $j\in\mathbb{N}$ we have $w^{(j)}\in{\mathcal{B}}_{m_{j}}$, and $w^{(j)}$ as a prefix of $w^{(j+1)}$. We will denote $\alpha_{j}=|w^{(j)}|$. We then define $w\in\mathscr{A}^{\infty}$ to be the unique point such that for every $j\in\mathbb{N}$ we have $w_{[0,\alpha_{j})}=w^{(j)}$.
Let $m_{1}\geq 4$ be such that $s(m_{1})>k_{1}$. By Lemma 6.11, there exists $w^{(1)}\in B_{m_{1}}$ so that
@@
The construction follows by taking $m_{j+1}>m_{j}$ satisfying $s(m_{j+1})>k_{j+1}$. Note that $m_{1}\geq 4$ implies $m_{j}\geq j+3$ for all $j\in\mathbb{N}$. By Lemma 6.12, there exists $w^{(j+1)}\in{\mathcal{B}}_{m_{j+1}}$ with $w^{(j)}$ as a prefix, and so that for every $\ell\in[\alpha_{j},\alpha_{j+1})$ it holds
@@
The minimal shift space $X$ has the asymptotic average shadowing property.
@@ 7 Applications
The equivalence of the vague specification and the asymptotic average shadowing properties allows us to use results already known about one property when dealing with the other. Consequently, we can address questions raised about systems exhibiting these properties.
Kamae showed in [11] that the specification property together with expansivity are sufficient conditions for the vague specification property. However, it is already known that the specification property alone implies the asymptotic average shadowing property (see [14, 16, 17]). Therefore, we can remove the assumption of expansivity from Kamae’s result.
@@
The rest of the section is devoted to answer questions from [7] and [14].
Question 7.2 ([7]).
@@
To answer this question we rely on results from [14] and [17]. We decided to skip the definitions of the weak specification and the almost specification property, and only invoke the results that are relevant to our purpose. We refer [16] for the definitions and further information.
Theorem 7.6 (Theorem 18 in [17]).
@@
The converse is not true and the almost specification property is used to prove it. Due to [14], it is known that the asymptotic average shadowing property (and consequently the vague specification property) is a consequence of the almost specification property.
Theorem 7.8 (Theorem 3.5 in [14]).
@@
It is known that the weak specification and the almost specification properties are independent of each other. Indeed, Theorem 17 in [18] shows that there are topological dynamical systems with almost specification but without weak specification. In Section 6 of [18] there is a construction of a topological dynamical system with weak specification, but without almost specification (we also refer to [14, 16, 18, 19] for more details).
@@
In Section 6.2 we present the construction of a class of minimal shift spaces with the $\bar{d}$-shadowing property from [6] (in particular, these shift spaces have the average shadowing property). By Corollary 6.15, every shift space in this class also has the asymptotic average shadowing property, which shows that Question 7.3 has a positive answer.
@@
The relation between the asymptotic average shadowing property and the almost specification is already clarified in [16].
A step toward an answer for the second part of Question 7.4 is given by Theorem 4.5 for surjective topological dynamical systems and by Corollary 5.17 for surjective shift spaces. Although we could not provide a definitive answer to the question, the lack of examples with the average shadowing property but not Besicovitch complete suggests a potential positive answer.
@@
Theorem 4.7 also provides a partial answer to this question. Indeed, we can use Proposition 2 in [11], which says that the vague specification property is inherited by factors, to confirm that the asymptotic average shadowing property is inherited by factors.
@@ Acknowledgements
Research of MEC was supported by NCN Polonez Bis grant no. 2021/43/P/ST1/02885 and NCN Polonez Bis grant No. 2022/47/P/ST1/00854; AT was supported by NCN Polonez Bis grant no. 2021/43/P/ST1/02885. The authors wish to express their gratitude to Dominik Kwietniak for suggesting the problem and providing insightful discussions.

### v2

@@ 6 Examples of shift spaces with the vague specification property
6.1 Proximal shift space with the vague specification property
We will show that the shift space $Z$, defined in Example 21 in [CKKK24], has the vague specification property. Since it is known that $Z$ has the $\bar{d}$-shadowing property, it is enough to show that $Z$ is $\bar{d}$-complete and use Corollary 5.17. Before showing that $Z$ is $\bar{d}$-complete we recall its construction following [CKKK24].
@@ 6.1 Proximal shift space with the vague specification property
We let $\tilde{E}_{j}$ and $E_{i}$ to be defined as in (10). Let $\{x^{(n)}\}_{n=1}^{\infty}$ be a $\bar{d}$-Cauchy sequence in $Z$. Passing to a subsequence, we can assume that for every $n\in\mathbb{N}$ the following holds
@@
Note that $Y_{n}$ is a hereditary shift with specification, in particular, synchronized, for every $n\in\mathbb{N}$. Therefore, by Lemmas 5.1 and 5.2, for every $n\in\mathbb{N}$ there exists $m_{n}\in\mathbb{N}$ such that $u_{n}=0^{m_{n}}$ is a synchronizing word for $Y_{n}$.
@@
Since $Z$ is hereditary, it follows that $y^{(n)}\in Z$ for every $n\in\mathbb{N}$. Observe that, using (11), for every $n\in\mathbb{N}$ we have
@@
The proximal shift space $Z$ has the vague specification property.
6.2 A class of minimal shift spaces with the vague specification property
In [CKKK24], the authors provided a way of constructing minimal shift spaces that are mixing and have the $\bar{d}$-shadowing property, acknowledging Piotr Oprocha for the ideas that inspired this construction. It is worth mentioning that the obtained shift spaces have positive topological entropy (for definition and details, we refer to [CKKK24]). We will show that those shift spaces are $\bar{d}$-complete.
@@ 6.2 A class of minimal shift spaces with the vague specification property
For $n\in\mathbb{N}$, let $X_{n}$ be the coded shift space generated by the code ${\mathcal{B}}_{n}$. That is, $X_{n}$ consists of all infinite concatenations of words from ${\mathcal{B}}_{n}$ together with their shifts. The proofs of the facts below can be found in Section 4.2 of [CKKK24].
@@
For every $n\in\mathbb{N}$ one has $X_{n+1}\subseteq X_{n}$, which implies $\bigcap_{n=1}^{\infty}X_{n}\neq\emptyset$.
@@
From now on, we set ${\mathcal{B}}_{1}=\{0,11\}$. As noted in [CKKK24], from this choice of ${\mathcal{B}}_{1}$, a simple inductive argument shows that, for every $n\in\mathbb{N}$, the set of lengths of all words in ${\mathcal{B}}_{n}$ is an interval.
@@
The next lemma is a direct consequence of Proposition 6.8 and was extracted from the proof of Proposition 25 in [CKKK24]. For the sake of completeness we will replicate the argument.
@@
In [CKKK24], it is proved that under certain conditions on the sequence $\{t(n)\}_{n=1}^{\infty}\subset\mathbb{N}$, the obtained shift space $X$ possesses the dynamical properties, as stated below.
Theorem 6.10 (Theorem 27 in [CKKK24]).
@@
The sequences $\{t(n)\}_{n=1}^{\infty}$ mentioned in Theorem 6.10 fulfill the following conditions for every $n\in\mathbb{N}$:
@@
From now on, we will assume that the sequence $\{t(n)\}_{n=1}^{\infty}$ satisfies the conditions above. We will show that under these conditions the shift space $X$ is $\bar{d}$-complete.
@@
We can improve Lemma 6.11 to ensure that the prefix of the obtained word $w$ can be predetermined, while also providing more precise estimates on the number of positions at which subwords of $x_{[0,|w|)}$ and $w$ differ.
@@
Let $\{x^{(n)}\}_{n=1}^{\infty}\subset X$ be a $\bar{d}$-Cauchy sequence. Passing, if necessary, to a subsequence we can assume that for every $n\in\mathbb{N}$ the following holds
@@
There exists a strictly increasing sequence $\{k_{n}\}_{n=1}^{\infty}\subset\mathbb{N}$ such that for every $n\in\mathbb{N}$ and $\ell\geq k_{n}$ we have
@@
We will construct a strictly increasing sequence $\{m_{j}\}_{j=1}^{\infty}\subset\mathbb{N}$ and a sequence of words $\{w^{(j)}\}_{j=1}^{\infty}\subset\mathcal{L}(X)$ such that for all $j\in\mathbb{N}$ we have $w^{(j)}\in{\mathcal{B}}_{m_{j}}$, and $w^{(j)}$ is a prefix of $w^{(j+1)}$. We will denote $\alpha_{j}=|w^{(j)}|$. We then define $w\in\mathscr{A}^{\infty}$ to be the unique point such that for every $j\in\mathbb{N}$ we have $w_{[0,\alpha_{j})}=w^{(j)}$.
Let $m_{1}\geq 4$ be such that $s(m_{1})>k_{1}$. By Lemma 6.11, there exists $w^{(1)}\in{\mathcal{B}}_{m_{1}}$ such that
@@
The construction follows by taking $m_{j+1}>m_{j}$ satisfying $s(m_{j+1})>k_{j+1}$. Note that $m_{1}\geq 4$ implies $m_{j}\geq j+3$ for all $j\in\mathbb{N}$. By Lemma 6.12, there exists $w^{(j+1)}\in{\mathcal{B}}_{m_{j+1}}$ with $w^{(j)}$ as a prefix, and so that for every $\ell\in[\alpha_{j},\alpha_{j+1})$ it holds that
@@
The minimal shift space $X$ has the vague specification property.
@@ 7 Applications
The equivalence of the vague specification and the asymptotic average shadowing properties allows us to use results already known about one property when dealing with the other. Consequently, we can address questions raised about systems that exhibit these properties.
Kamae showed in [Kam76] that the specification property together with expansivity are sufficient conditions for the vague specification property. However, it is already known that the specification property alone implies the asymptotic average shadowing property (see [KKO14, KŁO16, KŁO17]). Therefore, we can remove the assumption of expansivity from Kamae’s result.
@@
The rest of the section is devoted to answering questions from [DW24] and [KKO14].
Question 7.2 ([DW24]).
@@
To answer this question, we rely on results from [KKO14] and [KŁO17]. We decided to skip the definitions of the weak specification and the almost specification properties, and only invoke the results that are relevant to our purpose. We refer to [KŁO16] for definitions and further information.
Theorem 7.6 (Theorem 18 in [KŁO17]).
@@
The converse is not true, and the almost specification property is used to prove it. Due to [KKO14], it is known that the asymptotic average shadowing property (and consequently the vague specification property) is a consequence of the almost specification property.
Theorem 7.8 (Theorem 3.5 in [KKO14]).
@@
It is known that the weak specification and the almost specification properties are independent of each other. Indeed, Theorem 17 in [KOR16] shows that there are topological dynamical systems with almost specification but without weak specification. In Section 6 of [KOR16] there is a construction of a topological dynamical system with weak specification, but without almost specification (we also refer to [KKO14, KŁO16, KOR16, Pav16] for more details).
@@
In Section 6.2 we present the construction of a class of minimal shift spaces with the $\bar{d}$-shadowing property from [CKKK24] (in particular, these shift spaces have the average shadowing property). By Corollary 6.15, every shift space in this class also has the asymptotic average shadowing property, which shows that Question 7.3 has a positive answer.
@@
The relation between the asymptotic average shadowing property and the almost specification is already clarified in [KŁO16].
A step toward an answer for the second part of Question 7.4 is given by Theorem 4.4 for surjective topological dynamical systems and by Corollary 5.17 for surjective shift spaces. Although we were unable to provide a definitive answer to the question, the lack of examples with the average shadowing property but not Besicovitch complete suggests a potential positive answer.
@@
Theorem 4.6 also provides a partial answer to this question. Indeed, we can use Proposition 2 in [Kam76], which says that the vague specification property is inherited by factors, to confirm that the asymptotic average shadowing property is inherited by factors.
@@ Acknowledgements
The authors wish to express their gratitude to Dominik Kwietniak for suggesting the problem and providing insightful discussions. We also thank the anonymous reviewers for their careful reading and valuable comments that helped us improve this work.